OpenStudy (anonymous):

medal and fan! newbie to openstudy :) What is the distance from the coordinate pair (1, -3) to the coordinate pair (-3, 4)? Round your answer to the nearest hundredth. a. 3.31 b. 8.06 c. 7.93 d. 7.81

OpenStudy (phebe):

D or A but im going with A k

OpenStudy (anonymous):

are you sure? does anyone else have an opinion?

OpenStudy (phebe):

yes im sure

OpenStudy (anonymous):

We can't find distance until we know the slope so y2-y1/x2-x1=m(slope) and then use the equation of a line: y-y1=m(x-x1)

OpenStudy (anonymous):

I tried that, it gave me a completely wrong answer @jtryon

OpenStudy (anonymous):

@jtryon please please help :(

OpenStudy (anonymous):

ok... let me check your work

OpenStudy (anonymous):

so what did you get for slope?

OpenStudy (anonymous):

I did it on a piece of paper though

OpenStudy (anonymous):

slope = ?

OpenStudy (anonymous):

I don't know.. Should I go with A?

OpenStudy (anonymous):

I want you to understand the process

OpenStudy (anonymous):

slope which is m is equal to y2-y1/x2-x1 = -7/4

OpenStudy (anonymous):

oh i get it, I tried it again!!! so it is in fact A

OpenStudy (anonymous):

there you go :)

OpenStudy (anonymous):

thanks!!!